REU 2008: SCHEDULE FOR WEEK EIGHT

Tuesday will be a free day for mentors and mentees to meet and work on papers and practice talks.

All talks will be in E206. Lunches Monday, Wednesday, and Thursday will be in the tea room (spilling out into the hall). Lunch, dinner, and party Friday will be in the barn.

Wednesday at 2:15 we plan a question and answer open discussion session about graduate study in mathematics. All questions about what its like, what options are available, how best to prepare, what are the prospects, and so forth will be fielded.

We have three special faculty lectures, by Shmuel Weinberger, Jon Alperin, and Tathagata Basak in addition to talks by REU participants and mentors.

Times Monday Aug. 11 Wednesday Aug 13 Thursday Aug 14  Friday Aug 15
11:15 William Abram  Daniel Feltey  Noah Schweber 11:00 Evan Jenkins
12:00  lunch  lunch  lunch  lunch
1:00  Stephen Patrias Shmuel Weinberger  Nathan Schulz  Alex Rosenfeld
1:45  Alex Fix  followed by  Joseph Doliner  John Binder
2:30  Emily Riehl  Grad School Q and A Tathagata Basak  Jon Alperin
3:30  Tea  Tea  Tea  Tea
4:00  Elan Bechor  Sarah Bennett  Caitlin Waddle  Carmel Levy
4:45 Damir Dzhafarov  Vipul Naik  Ian Biringer  Mike Shulman
6:00  ——–  ——–  ——–  Dinner/Party